Consider the system of equations

2x1+x2+4x3=11

3x1+x2+5x3=14

feasible solution is x1=2,x2=3,x3 =1.Reduce this feasible solution to a basic feasible solution.



1
Expert's answer
2021-06-21T12:38:26-0400

"2x_1+x_2=11-4x_3"

"3x_1+x_2=14-5x_3"


"\Delta=\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}=-1"


"\Delta_1=\begin{vmatrix}
 11-4x_3 & 1 \\
 14-5x_3 & 1
\end{vmatrix}=x_3-3"


"\Delta_2=\begin{vmatrix}
 2 & 11-4x_3 \\
 3 & 14-5x_3
\end{vmatrix}=2x_3-5"


"x_1=\Delta_1/\Delta=3-x_3"


"x_2=\Delta_2/\Delta=5-2x_3"